# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v2

@@ II-B Distributed Dual-Gradient Method
The Lagrangian function corresponding to the optimization problem in (4) is given by
($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda)$ $\!=\!\sum_{i=1}^{M}\!J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})\!+\!\lambda^{\top}\left(\sum_{i=1}^{M}f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-b(\varepsilon)\!\right)$ (10)
$=\sum_{i=1}^{M}\left(J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+\lambda^{\top}g_{i}(\tilde{{\bm{u}}}_{i})\right),$
where $\lambda\in\mathbb{R}^{Np}_{+}$ (the non-negative orthant of $\mathbb{R}^{Np}$) is the Lagrangian multiplier and $g_{i}(\tilde{{\bm{u}}}_{i}):=f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-\frac{b(\varepsilon)}{M}$. The dual problem of (4) is defined as
@@
Based on Assumptions 1, 2 and the definition of DMPC problem, it can be concluded that the strong duality holds for (4), and the optimization problem (4) can be addressed by solving its dual problem (11). In addition, the Saddle-Point Theorem holds, i.e., given an optimal primal-dual pair ($(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$, the following relationship holds for any $\lambda\in\mathbb{R}^{Np}_{+}$ and $\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$:
($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda)\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda^{*}).$ (12)
A conventional approach to solving problem (11) is the distributed dual-gradient method [27, 28]. The core idea is to regard the Lagrangian multiplier (dual variable) $\lambda$ as a consensus variable and then $M$ subsystems address the optimization problem in a collective manner. Specifically, let each subsystem have a local copy $\lambda_{i}^{k}$ of the dual variable. $\Pi_{\mathbb{R}^{Np}_{+}}\left[\cdot\right]$ denotes Euclidean projection of a vector on the set +$\mathbb{R}^{Np}_{+}$, and $\gamma^{k}>0$ denotes the step-size. Then, the distributed dual-gradient method is summarized in Algorithm 1, and the overall DMPC implementation is detailed in Algorithm 2.
@@ 4             Every subsystem $i$ sends $\lambda_{i}^{k}$ to subsystem $j\in\mathcal{N}_{i}$;
1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
2Every subsystem $i$ computes $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 1 with $x_{i}(t)$;
@@ 1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
If Assumptions 2 and 3 hold and if the step-size $\gamma^{k}$ satisfies the conditions $\sum_{k=0}^{\infty}\gamma^{k}=\infty$, $\sum_{k=0}^{\infty}(\gamma^{k})^{2}<\infty$, then Algorithm 1 guarantees the convergence of the sequence $\{\tilde{{\bm{u}}}_{i}^{k}\}$, i.e., $\lim_{k\rightarrow\infty}\|\tilde{{\bm{u}}}_{i}^{k}-\tilde{{\bm{u}}}_{i}^{*}\|$. Note that the objective function of problem (4) is strictly convex, and thus the asymptotic primal convergence can be established without resorting to local averaging mechanism (see Section 3.4 in [28] for more details). In addition, if the iteration number $\bar{k}$ is selected sufficiently large to meet specific criteria [29, 11], then Algorithm 2 ensures the feasibility and stability of the considered MPC problem.
In Algorithm 1, each subsystem can avoid sharing the primal variable and only share its local copy $\lambda_{i}^{k}$ of the dual variable with its neighbors. However, this sharing mechanism cannot provide strong privacy protection as the iteration trajectory of $\lambda_{i}^{k}$ still bears information of the primal variable. In particular, we assume that the adversary has prior knowledge about the communication network $L$ and the step-size $\gamma^{k}$, and can get access to all information exchanged in communication channels. Under this circumstance, the adversary can record the updates of $\bar{\lambda}_{i}^{k}$ and $\lambda_{i}^{k}$ at each iteration. Then, based on $\bar{\lambda}_{i}^{k}$ and $\lambda_{i}^{k+1}$ in two consecutive iterations and $\gamma^{k}$, the adversary can employ (15) to estimate the value of $g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})$. It should be noted that $g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})$ is privacy-sensitive as it is the function of primal variable and is used to formulate the coupled global constraint. Therefore, it is necessary to incorporate a privacy protection mechanism into the distributed dual-gradient algorithm such that the DMPC problem can be addressed with privacy protection.
@@ II-C On Differential Privacy
In this work, DP is used to characterize and quantify the achieved privacy level of distributed optimization algorithms. Given the continual exchange of information among subsystems in iterative optimization algorithms, the notion of $\epsilon$-DP for continuous bit streams [30] is adopted. Drawing inspiration from the distributed optimization framework proposed by [22], we represent the DMPC problem in (4) by four parameters $(L,\mathcal{J},\tilde{\mathcal{U}},\mathcal{G})$ to facilitate DP analysis. Specifically, $L$ is the interaction weight matrix describing the communication network, $\mathcal{J}:=\{J_{1},\,\cdots,J_{M}\}$ denotes the set of objective functions for individual subsystems, $\tilde{\mathcal{U}}:=\{\tilde{\mathcal{U}_{1}},\cdots,\tilde{\mathcal{U}_{M}}\}$ is the domain of optimization variables, and $\mathcal{G}:=\{g_{1},\cdots,g_{M}\}$ represents the set of constraint functions for individual subsystems. The adjacency between two optimization problems is defined as follows:
@@
Two distributed optimization problems $\mathcal{P}=(L,\mathcal{J},\tilde{\mathcal{U}},\mathcal{G})$ and $\mathcal{P}^{{}^{\prime}}=(L^{{}^{\prime}},\mathcal{J}^{{}^{\prime}},\tilde{\mathcal{U}}^{{}^{\prime}},\mathcal{G}^{{}^{\prime}})$ are adjacent if they satisfy the following conditions:
  • the interaction weight matrices, the objective functions, and the domains of optimization variables are identical, i.e., $L=L^{{}^{\prime}}$, $\mathcal{J}=\mathcal{J}^{{}^{\prime}}$, and $\tilde{\mathcal{U}}=\tilde{\mathcal{U}}^{{}^{\prime}}$;
  • there exists an $i\in\mathbb{Z}_{1}^{M}$ such that $g_{i}\neq g_{i}^{{}^{\prime}}$, and $g_{j}=g_{j}^{{}^{\prime}}$ for all $j\in\mathbb{Z}_{1}^{M},\,j\neq i$;
  • $g_{i}$ and $g_{i}^{{}^{\prime}}$, while different, exhibit similar behaviors near $\theta^{*}$, where $\theta^{*}$ is the solution of $\mathcal{P}$. More precisely, there exists a $\delta>0$ such that for all $\bm{u}_{i}$ and $\bm{u}_{i}^{{}^{\prime}}$ within the domain $B_{\delta}(\theta^{*}):=\{\bm{v}:\bm{v}\in\mathbb{R}^{Nm_{i}},\|\bm{v}-\theta^{*}\|<\delta\}$, $g_{i}(\bm{u}_{i})=g_{i}^{{}^{\prime}}(\bm{u}_{i}^{{}^{\prime}})$ holds.
We denote the execution of a distributed optimization algorithm as $\mathcal{A}$, represented by a sequence of the iteration variable $\vartheta$, i.e., $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\cdots\}$. We assume that adversaries have access to all communicated messages among the subsystems. Hence, under an execution $\mathcal{A}$, the adversaries’ observation is the sequence of communicated messages, denoted as $\mathcal{O}$. Let $\mathbb{O}$ represent the set of all possible observation sequences. For a distributed optimization problem $\mathcal{P}$ with an initial state $\vartheta^{0}$, the observation mapping is defined as $\mathcal{R}_{\mathcal{P},\vartheta^{0}}(\mathcal{A}):=\mathcal{O}$. Furthermore, for a distributed optimization problem $\mathcal{P}$, an initial state $\vartheta^{0}$, and an observation sequence $\mathcal{O}$, $\mathcal{R}_{\mathcal{P},\vartheta^{0}}^{-1}(\mathcal{O})$ denotes the set of executions $\mathcal{A}$ capable of generating the observation $\mathcal{O}$.
Definition 2 ($\epsilon$-differential privacy, [22]).
@@
The definition of $\epsilon$-DP guarantees that adversaries, with access to all communicated information, cannot infer knowledge about any participating subsystem’s sensitive information. It can be found that a smaller $\epsilon$ indicates a better extent of privacy preservation.
@@ III-A Algorithm Description
In this section, a DP noise injection mechanism is proposed to achieve privacy preservation in the distributed dual-gradient algorithm. The developed algorithm is summarized in Algorithm 3.
In contrast to Algorithm 1, where each subsystem directly sends $\lambda_{i}^{k}$ to its neighbors, Algorithm 3 incorporates DP noise $\zeta_{i}^{k}$ into $\lambda_{i}^{k}$ and shares the perturbed signal $\hat{\lambda}_{i}^{k}:=\lambda_{i}^{k}+\zeta_{i}^{k}$ among the communication network. Therefore, the information available to potential adversaries is the sequence $\{\hat{\lambda}_{i}^{k}\}$. Due to the randomness of DP noise, it is impossible for the adversary to extract useful information from $\{\hat{\lambda}_{i}^{k}\}$ with significant probability. Furthermore, it should be noted that directly integrating persistent DP noise into existing optimization algorithms will compromise the convergence accuracy. To address this issue, we utilize findings from [25, 26] to design a weakening factor. As shown in (18), the weakening factor, denoted as $\chi^{k}$, is applied on the interaction terms ($L_{ij}(\hat{\lambda}_{j}^{k}-\lambda_{i}^{k})$). The fundamental principle behind incorporating this weakening factor is to gradually eliminate the impact of DP noise on convergence accuracy.
@@
[…]
@@
where $\{\chi^{k}\}$ is the weakening factor sequence from Algorithm 3.
@@
In Algorithm 3, we allow the variance of DP noise $\zeta_{i}^{k}$, i.e., $2(\nu^{k})^{2}$, to be constant or increasing with $k$. To satisfy condition (19), one can carefully design the weakening factor sequence $\{\chi^{k}\}$ to make its decreasing rate outweigh the increasing rate of the noise level sequence $\{\nu^{k}\}$. For instant, (19) can be satisfied by setting $\chi^{k}=\frac{c_{1}}{1+c_{2}k^{c_{3}}}$ and $\nu^{k}=d_{1}+d_{2}k^{d_{3}}$ with any $c_{1}>0$, $c_{2}>0$, $10.5<c_{3}<$ 1, $d_{1}>0$, $d_{2}>0$, and $0<d_{3}<0.5-c_{3}$.
@@ III-B Convergence Analysis
Suppose Assumptions 1, 3, and 4 hold. If the non-negative weakening factor sequence $\{\chi^{k}\}$ and the step-size sequence $\{\gamma^{k}\}$ in Algorithm 3 satisfy
$\sum_{k=0}^{\infty}\chi^{k}=\infty,\,\sum_{k=0}^{\infty}(\chi^{k})^{2}<\infty,\,\sum_{k=0}^{\infty}\frac{(\gamma^{k})^{2}}{\chi^{k}}<\infty,$ (21)
then the following results hold almost surely:
  1. 1.
    $\lim_{k\to\infty}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|=0$ for all $i\in\mathbb{Z}_{1}^{M}$;
  2. 2.
    $\sum_{k=0}^{\infty}\chi^{k}\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|^{2}<\infty$;
  3. 3.
    $\sum_{k=0}^{\infty}\gamma^{k}\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|<\infty$.
@@
Based on Assumption 1 and (7), it can be concluded that the local constraint set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$ is bounded. Then, from (8) and the relation $g_{i}(\tilde{{\bm{u}}}_{i}):=f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-\frac{b(\varepsilon)}{M}$, we have that for any $\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$, $g_{i}(\tilde{{\bm{u}}}_{i})$ is bounded, i.e., there exists a constant $C_{g}\in\mathbb{R}_{+}$ such that
$\|g_{i}(\tilde{{\bm{u}}}_{i})\|\leq C_{g},\forall\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t)),i\in\mathbb{Z}_{1}^{M}.$ (22)
According to Assumptions 3, 4, (21), and (22), we can follow the same line of reasoning as that of Theorem 1 in [26] to obtain the results. ∎
The following lemma is also required for convergence analysis:
Lemma 1 (Lemma 11, [31]).
Let $\{\psi^{k}\}$, $\{\phi^{k}\}$, $\{a^{k}\}$, and $\{b^{k}\}$ be random non-negative scalar sequences such that
$\mathbb{E}\left[\psi^{k+1}|\mathcal{F}^{k}\right]\leq(1+{a^{k}})\psi^{k}-\phi^{k}+{b^{k}},\quad\forall k\geq 0,$
[…]
@@
Suppose Assumptions 1, 3, and 4 hold. If the non-negative sequences $\{\chi^{k}\}$ and $\{\gamma^{k}\}$ satisfy $\sum_{k=0}^{\infty}\chi^{k}=\infty$, $\sum_{k=0}^{\infty}(\chi^{k})^{2}<\infty$, $\sum_{k=0}^{\infty}\gamma^{k}=\infty$, and $\sum_{k=0}^{\infty}\frac{(\gamma^{k})^{2}}{\chi^{k}}<\infty$, then Algorithm 3 guarantees that
$\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*}),$
$\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k}\},\lambda^{*})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$
hold almost surely.

### v3

@@ II-B Distributed Dual-Gradient Method
The Lagrangian function corresponding to the optimization problem in (3) is given by $\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda)=\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+\lambda^{\top}\left(\sum_{i=1}^{M}f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-b(\varepsilon)\right)=\sum_{i=1}^{M}\left(J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+\lambda^{\top}g_{i}(\tilde{{\bm{u}}}_{i})\right)$, where $\lambda\in\mathbb{R}^{Np}_{+}$ is the Lagrangian multiplier and $g_{i}(\tilde{{\bm{u}}}_{i}):=f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-\frac{b(\varepsilon)}{M}$. The dual problem of (3) is defined as
@@
Under Assumptions 1, 2, strong duality holds for (3), and the optimization problem (3) can be solved via its dual formulation (9). In addition, the Saddle-Point Theorem holds, i.e., given an optimal primal-dual pair $(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$, the following relationship holds for any $\lambda\in\mathbb{R}^{Np}_{+}$ and $\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$:
$\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda)\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda^{*}).$ (10)
A standard approach for solving (9) is the distributed dual-gradient method [Falsone2017AUTO, Notarstefano2019FT]. Specifically, the dual variable $\lambda$ is treated as a consensus variable, and each subsystem has a local copy $\lambda_{i}^{k}$. $\Pi_{\mathbb{R}^{Np}_{+}}\left[\cdot\right]$ denotes Euclidean projection of a vector on the set $\mathbb{R}^{Np}_{+}$, and $\gamma^{k}>0$ is the step-size. Then, the distributed dual-gradient method is summarized in Algorithm 1, and the overall DMPC implementation is detailed in Algorithm 2.
In Algorithm 1, each subsystem avoids sharing the primal variable and only shares its local copy $\lambda_{i}^{k}$ of the dual variable with its neighbors. However, this sharing mechanism cannot provide strong privacy protection, as the iteration trajectory of $\lambda_{i}^{k}$ still bears information of the primal variable. In particular, since the communication network $L$ and the step-size $\gamma^{k}$ are public information (otherwise the algorithm cannot be implemented in a fully decentralized manner), if an adversary can intercept all information exchanged in communication channels, it can record the updates of $\tilde{\lambda}_{i}^{k}$ and $\lambda_{i}^{k}$ at each iteration. Using consecutive updates $\tilde{\lambda}_{i}^{k}$ and $\lambda_{i}^{k+1}$, along with $\gamma^{k}$, the adversary can use (13) to estimate $g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})$. The value of $g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})$ is privacy-sensitive, as it depends on the primal variable and is used to formulate the coupled global constraint. Therefore, it is necessary to incorporate a privacy protection mechanism into the distributed dual-gradient algorithm to ensure rigorous privacy protection in DMPC.
@@ 4     Every subsystem $i$ sends $\lambda_{i}^{k}$ to subsystem $j\in\mathcal{N}_{i}$;
1At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
2Every subsystem $i$ computes $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 1;
@@ II-C On Differential Privacy
In this work, DP is used to characterize and quantify the achieved privacy level of distributed optimization algorithms. Drawing inspiration from the distributed optimization framework proposed by [Huang2015], we represent the DMPC problem in (3) by four parameters $(L,\mathcal{J},\tilde{\mathcal{U}},\mathcal{G})$ to facilitate DP analysis. Specifically, $L$ is the interaction weight matrix describing the communication network, $\mathcal{J}:=\{J_{1},\,\cdots,J_{M}\}$ denotes the set of objective functions for individual subsystems, $\tilde{\mathcal{U}}:=\{\tilde{\mathcal{U}_{1}},\cdots,\tilde{\mathcal{U}_{M}}\}$ is the domain of optimization variables, and $\mathcal{G}:=\{g_{1},\cdots,g_{M}\}$ represents the set of constraint functions for individual subsystems. The adjacency between two optimization problems is defined as follows:
@@
Two distributed optimization problems $\mathcal{P}=(L,\mathcal{J},\tilde{\mathcal{U}},\mathcal{G})$ and $\mathcal{P}^{{}^{\prime}}=(L^{{}^{\prime}},\mathcal{J}^{{}^{\prime}},\tilde{\mathcal{U}}^{{}^{\prime}},\mathcal{G}^{{}^{\prime}})$ are adjacent if they satisfy the following conditions: $\left.1\right)$ $L=L^{{}^{\prime}}$, $\mathcal{J}=\mathcal{J}^{{}^{\prime}}$, and $\tilde{\mathcal{U}}=\tilde{\mathcal{U}}^{{}^{\prime}}$; $\left.2\right)$ There exists an $i\in\mathbb{Z}_{1}^{M}$ such that $g_{i}\neq g_{i}^{{}^{\prime}}$, and $g_{j}=g_{j}^{{}^{\prime}}$ for all $j\in\mathbb{Z}_{1}^{M},\,j\neq i$; $\left.3\right)$ $g_{i}$ and $g_{i}^{{}^{\prime}}$, while different, exhibit similar behaviors near $\theta^{*}$, where $\theta^{*}$ is the solution of $\mathcal{P}$. More precisely, there exists a $\delta>0$ such that for all $\bm{u}_{i}$ and $\bm{u}_{i}^{{}^{\prime}}$ within the domain $B_{\delta}(\theta^{*}):=\{\bm{v}:\bm{v}\in\mathbb{R}^{Nm_{i}},\|\bm{v}-\theta^{*}\|<\delta\}$, $g_{i}(\bm{u}_{i})=g_{i}^{{}^{\prime}}(\bm{u}_{i}^{{}^{\prime}})$ holds.
We denote the execution of a distributed optimization algorithm as $\mathcal{A}$, represented by a sequence of the iteration variable $\vartheta$, i.e., $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\cdots\}$. Assuming adversaries have access to all communicated messages among subsystems, their observation under an execution $\mathcal{A}$ is the sequence of these messages, denoted by $\mathcal{O}$. Let $\mathbb{O}$ represent the set of all possible observation sequences. For a distributed optimization problem $\mathcal{P}$ with an initial state $\vartheta^{0}$, the observation mapping is defined as $\mathcal{R}_{\mathcal{P},\vartheta^{0}}(\mathcal{A}):=\mathcal{O}$. Moreover, given $\mathcal{P}$, $\vartheta^{0}$, and an observation sequence $\mathcal{O}$, $\mathcal{R}_{\mathcal{P},\vartheta^{0}}^{-1}(\mathcal{O})$ denotes the set of executions $\mathcal{A}$ that could generate the observation $\mathcal{O}$.
Definition 2 ($\epsilon$-differential privacy, [Huang2015]).
@@
The definition of $\epsilon$-DP guarantees that adversaries, with access to all communicated information, cannot infer knowledge about any participating subsystem’s sensitive information.
@@ III-A Algorithm Description
In this section, a DP noise injection mechanism is proposed to achieve privacy preservation in the distributed dual-gradient algorithm. The method is summarized in Algorithm 3.
In contrast to Algorithm 1, where each subsystem directly sends $\lambda_{i}^{k}$ to its neighbors, Algorithm 3 adds DP noise $\zeta_{i}^{k}$ to $\lambda_{i}^{k}$ and shares the perturbed signal $\hat{\lambda}_{i}^{k}\!:=\!\lambda_{i}^{k}\!+\!\zeta_{i}^{k}$ among the communication network. Thus, the adversary’s available information is the sequence $\{\hat{\lambda}_{i}^{k}\}$. The randomness introduced by the DP noise ensures that extracting meaningful information from $\{\hat{\lambda}_{i}^{k}\}$ is statistically impossible. Moreover, it should be noted that directly integrating persistent DP noise into optimization algorithms will compromise the convergence accuracy. To address this issue, we utilize findings from [WangTAC2024, WangTAC2023] to design diminishing weakening factor sequence $\{\chi^{k}\}$ and step-size sequence $\{\gamma^{k}\}$. As shown in (16), $\chi^{k}$ and $\gamma^{k}$ are applied on the terms ($L_{ij}(\hat{\lambda}_{j}^{k}\!-\!\lambda_{i}^{k})$) and $g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})$, respectively. The principle behind incorporating the diminishing weakening factor and step-size sequences is to gradually eliminate the impact of DP noise, thereby ensuring convergence accuracy.
@@
$\begin{aligned} \lambda_{i}^{k+1}=\Pi_{\mathbb{R}^{Np}_{+}}[\lambda_{i}^{k}&+\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\hat{\lambda}_{j}^{k}-\lambda_{i}^{k})\\ &+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})];\end{aligned}$ (16)
@@
where $\{\chi^{k}\}$ ($\chi^{k}>0$) is the weakening factor sequence from Algorithm 3.
@@
In Algorithm 3, the variance of DP noise $\zeta_{i}^{k}$, i.e., $2(\nu^{k})^{2}$, can be constant or increasing with $k$. To satisfy condition (17), one can carefully design the weakening factor sequence $\{\chi^{k}\}$ to make its decreasing rate outweigh the increasing rate of the noise level sequence $\{\nu^{k}\}$. For instant, (17) is satisfied with $\chi^{k}=\frac{c_{1}}{1+c_{2}k^{c_{3}}}$ and $\nu^{k}=d_{1}+d_{2}k^{d_{3}}$, where $c_{1}>0$, $c_{2}>0$, $0.5<c_{3}<1$, $d_{1}>0$, $d_{2}>0$, and $0<d_{3}\leq 1-c_{3}$. For notation simplicity, we assume that all subsystems use the same Laplace distribution ${\rm Lap}(\nu^{k})$ to generate DP noise. Actually, each subsystem can independently select its DP noise intensity as long as condition (17) is met.
@@ III-B Convergence Analysis
Suppose Assumptions 1, 3, and 4 hold. If the non-negative weakening factor sequence $\{\chi^{k}\}$ and the step-size sequence $\{\gamma^{k}\}$ in Algorithm 3 satisfy $\sum_{k=0}^{\infty}\chi^{k}=\infty$, $\sum_{k=0}^{\infty}(\chi^{k})^{2}<\infty$, and $\sum_{k=0}^{\infty}\frac{(\gamma^{k})^{2}}{\chi^{k}}<\infty$, then the following results hold almost surely: $\left.1\right)$ $\lim_{k\to\infty}\|\lambda_{i}^{k}-\bar{\lambda}^{k}\|=0$ for all $i\in\mathbb{Z}_{1}^{M}$; $\left.2\right)$ $\sum_{k=0}^{\infty}\chi^{k}\sum_{i=1}^{M}\|\lambda_{i}^{k}-\bar{\lambda}^{k}\|^{2}<\infty$; $\left.3\right)$ $\sum_{k=0}^{\infty}\gamma^{k}\sum_{i=1}^{M}\|\lambda_{i}^{k}-\bar{\lambda}^{k}\|<\infty$.
@@
As stated in Assumption 1, $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ are bounded, and then it can be concluded from (6) that the local constraint set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$ is bounded. From (7) and the relation $g_{i}(\tilde{{\bm{u}}}_{i}):=f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-\frac{b(\varepsilon)}{M}$, we have that for any $\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$, $g_{i}(\tilde{{\bm{u}}}_{i})$ is bounded, i.e., there exists a constant $C_{g}\in\mathbb{R}_{+}$ such that $\|g_{i}(\tilde{{\bm{u}}}_{i})\|\leq C_{g},\forall\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t)),i\in\mathbb{Z}_{1}^{M}$. According to Assumptions 3, 4, and the conditions that $\sum_{k=0}^{\infty}\chi^{k}=\infty$, $\sum_{k=0}^{\infty}(\chi^{k})^{2}<\infty$, $\sum_{k=0}^{\infty}\frac{(\gamma^{k})^{2}}{\chi^{k}}<\infty$, and $\|g_{i}(\tilde{{\bm{u}}}_{i})\|\leq C_{g}$, we can follow the same line of reasoning as that of Theorem 1 in [WangTAC2024] to obtain the results. ∎
The following lemma is required for convergence analysis:
Lemma 1 (Lemma 11, [Polyak1987]).
Let $\{\psi^{k}\}$, $\{\phi^{k}\}$, $\{a^{k}\}$, and $\{\varpi^{k}\}$ be random non-negative scalar sequences such that
$\mathbb{E}\left[\psi^{k+1}|\mathcal{F}^{k}\right]\leq(1+{a^{k}})\psi^{k}-\phi^{k}+{\varpi^{k}},\quad\forall k\geq 0,$
where $\mathcal{F}^{k}=\{\psi^{\ell},\phi^{\ell},{a^{\ell}},{\varpi^{\ell}};\,0\leq\ell\leq k\}$. If $\sum_{k=0}^{\infty}{a^{k}}<\infty$ and $\sum_{k=0}^{\infty}{\varpi^{k}}<\infty$, then $\sum_{k=0}^{\infty}\phi^{k}<\infty$ and $\{\psi^{k}\}$ converges to a finite variable almost surely.
@@
Suppose Assumptions 1, 3, and 4 hold. If the non-negative sequences $\{\chi^{k}\}$ and $\{\gamma^{k}\}$ satisfy $\sum_{k=0}^{\infty}\chi^{k}=\infty$, $\sum_{k=0}^{\infty}(\chi^{k})^{2}<\infty$, $\sum_{k=0}^{\infty}\gamma^{k}=\infty$, and $\sum_{k=0}^{\infty}\frac{(\gamma^{k})^{2}}{\chi^{k}}<\infty$, then Algorithm 3 guarantees that $\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ and $\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k}\},\lambda^{*})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ hold almost surely.